one number is 12 less than the other, their product is 288. find the greater number?
Answer by stanbon(75887)

one number is 12 less than the other, their product is 288.
find the greater number?
------
1st: x
2nd: x-12
-------
Equation:
x(x-12) = 288
x^2 - 12x - 288 = 0
-----
(x-24)(x+12) = 0
---
X = 24
x-12 = 12
=================
Cheers,
Stan H.